Suppose that A, B, CER55 and that |A| = -2, \B = 4, CI=1. Find the determinant of the following: 2A’B-2CT
Answer #1

Suppose A, B, C are matrices LA = -2 (3) = 4 - 1B11= 14 el =T ICT-T determinant of 2A3B² CT 2 1413 B-11² ICTI 2626 60 26-3)